A uniform ring of mass M and radius R is placed
directly above a uniform sphere of mass 8 M and of
same radius R. The centre of the ring is at a distance
of d = 3 R from the centre of the sphere. The
gravitational attraction between the sphere and the
ring is :

Grade:12

4 Answers

ramendra bahadur singh
13 years ago

F=G*8M2/16R2    OR    F=G*M2/2R2

Rajkuimar Yadav
13 years ago

ans. 24GM2/10(10)1/2R2

sourabh singh
13 years ago

treat sphere is a ppoint mass  and it is same as ring and a particle ...hope u rather know it to found it now

Manish
5 years ago
Gravitational field due to ring at a distanced=3–√R on its axis isE=GMd(R2+d2)3/2=3–√GM8R2Force on the sphere = mass of sphere × E=8ME=3–√GM2R2